solve each quadratic in form equation

1. 4y^4+9=13y^2
1
Expert's answer
2012-12-10T11:05:50-0500
let t=y^2 then
4t^2-13t+9=0
D=13^2-4*9*4=169-144=25
t12=(13+-5)/8 so t1=1, t2=9/4
but t=y^2
so y^2=1 or y^2=9/4
y1=1, y2=-1, y3=-3/2, y4=3/2